OpenStudy (anonymous):

What is the arc length when Θ = 4 pi / 7 and the radius is 5 cm? 20 pi over 7cm 4 pi over 7cm 2 pi over 7cm 24 pi over 7cm

OpenStudy (jdoe0001):

\(\bf \textit{arc length}=s=\theta\cdot r\)

OpenStudy (jdoe0001):

r = radius of course

OpenStudy (anonymous):

What does s stand for? @jdoe0001

OpenStudy (jdoe0001):

s = arc length

OpenStudy (jdoe0001):

\(\bf \textit{arc length}=s=\theta\cdot r\qquad r=5\qquad\theta=\cfrac{4\pi}{7}\qquad thus\qquad s=\cfrac{4\pi}{7}\cdot 5\)